# Optimizing reaction and transport fluxes in temperature gradient-driven chemical reaction-diffusion systems

### v1

@@ Abstract
We present a theoretical study of 1D chemical systems subjected to temperature gradients. A complete theoretical framework describes the system behavior induced by various temperature profiles. An exact mathematical derivation was established for a simple two-compartment model, and generalized to arbitrary reaction-diffusion systems based on numerical models. An experimental system was eventually scaled and tuned to optimize either nonequilibrium chemical transport or reaction.
The relevant parameters for this description were identified; they focused on the system symmetry for chemical reaction and transport. Nonequilibrium thermodynamic approaches lead to a description analogous to electric circuits. Temperature gradients lead to the onset of a steady chemical force, sustaining steady reaction-diffusion fluxes moderated by chemical resistance. The system activity was then assessed using the entropy production rate, as a measure of its dissipated power.
The chemical characteristics of the system can be tuned for the general optimization of the nonequilibrium state or for the specific optimization of either transport or reaction processes. The temperature gradient shape can be tailored to precisely control the spatial localization of active processes, targeting either precise spatial localization or propagation over large areas. The resulting temperature-driven chemical system can in turn be used to drive secondary processes into either nonequilibrium reaction fluxes or concentration gradients.
Keywords: Free energy transduction, coupled reactions, reaction-diffusion, steady-state, temperature gradient, thermodynamics of nonequilibrium, chemical fluxes, entropy rate.
preprint: Submitted to Journal of Chemical Physics
@@ I Introduction
Thermodynamics fixes the final composition at which a reactive system reaches equilibrium under chemical and physical constraints. In practice, it governs the realization of multiple chemical reactions of academic and industrial interest, as well as the elementary steps of most separation processes, such as distillation or chromatography. In contrast, kinetics drives the trajectory of the evolution of a reactive system. It also governs the dynamic behavior of the latter when sustained in an out-of-equilibrium steady-state, which can exhibit much richer phenomena (e.g. oscillations, waves, patterns, and chaos) than in the equilibrium state. In particular, such out-of-equilibrium steady-states are encountered in living biological matter, where they control its dynamics[1, 2] and lead to the propagation of a sustained energy flux throughout metabolism by exploiting free energy transduction.[3, 4, 5]
@@
The dynamic behavior of the out-of-equilibrium steady-states is theoretically understood.[6, 7, 8] In contrast, the issue of steadily sustaining an out-of-equilibrium reactive state has received much less attention and remains a major challenge.[9] In practice, its realization requires either introducing constraints to prevent the effective relaxation of the reactive processes towards equilibrium,[10, 11, 12] or steadily applying a periodic excitation on the reactive system.[13] We previously demonstrated the relevance of the first approach by frustrating the relaxation of reactive processes by the interference of molecular motion,[14, 15] a strategy to generate reaction-diffusion cycles that are thought to establish spatial gradients of signaling activities in living cells.[16, 17] However, the efficiency of this first approach has not been theoretically analyzed, and is the object of the present manuscript.
In our previous work,[14, 15] we used light as the energy source to sustain a reactive system in an out-of-equilibrium state. In the present study, we favored a more general approach and adopted the application of a temperature gradient. Despite its chemical relevance, the behavior of reactive media in contact with two thermostats at different temperatures has involved glorious names (including Nernst,[18] Dirac,[19] Prigogine[20, 21]). A few papers have been published since, but have been mostly concerned with the gas phase and the question of how a chemical reaction running in a system modifies its thermal conductivity.[22, 23, 24, 25, 26, 27, 28, 29, 30] Recently, heat flow has been applied to generate proton gradients and pH oscillations in microscale aqueous solutions.[31, 32] In the latter works, the reactive system was often assumed to be at chemical equilibrium within the temperature gradient.
@@
In this manuscript, we level off this assumption, which has been predicted to generate attractive behavior such as dissipation-driven selection of states in nonequilibrium chemical networks,[33] or emergent thermophoretic behavior in chemical reaction systems.[34] More specifically, we identify the optimal regime for shifting away a reactive system from chemical equilibrium at a steady-state. We analyzed the optimization of the flux of the reaction and diffusion processes and of the corresponding entropy production,[35] which are key elements in propagating energy into a set of coupled reactions. To achieve this optimization, we benefit from a fruitful analogy between entropy production by a chemical reaction and the ohm dissipated power by electrical resistors. We then identify the conditions for the spatial localization or delocalization of nonequilibrium activity. Finally, we assessed the possibility of propagating this primary energy transfer from the heat sources and sinks to secondary athermic reaction or transport processes.
@@ II.1 Chemical model
This represents a generic reaction, in its simplest form. It can be easily extended to more complex mechanisms, as detailed further, namely for the description of second order reactions or more complex mechanisms.
@@
Figure 1: General Scheme
@@ II.2 Chemical response to a temperature gradient between two homogeneous compartments
The general behavior of a temperature gradient-driven reaction was first characterized in the limit of two compartments, each of which was homogeneous in temperature and concentrations. This corresponds to two stirred systems; each compartment $A/B$ is connected to the other by chemical exchange of the reaction rate constants $\delta_{i}$ for each compound of concentration $u_{i}$.
This setup can be reduced to a set of ordinary differential equations, with one set of parameters for each compartment. Analytical solutions for the resulting steady-state were obtained. A second order Taylor development as a function of the temperature gradient $\theta_{g}$ leads to simpler—but more informative—solutions.
@@ II.2.1 Parameters
The identification of the relevant parameters for the description of the system is a critical step. We focused on describing the system in general terms by assessing the global symmetry or asymmetry of the parameters, as this leads to a description of the system behavior in the simplest mathematical terms.
The general properties of the system are described based on its characteristic dimensions: concentration $c_{0}$ (corresponding to the total concentration), temperature $T_{0}$ (corresponding to the median system temperature), and time $t_{0}$. All subsequent system parameters were nondimensionalized using these characteristic parameters (see Table 1).
The system is thus described in the nondimensionalized concentrations $u_{i,j}$ of compound $i$ in compartment $j$, relative temperature deviation $\theta$, and time $\bar{t}$. Each compartment is thermostated at temperature $\theta_{a}=-\theta_{g}$ and $\theta_{b}=+\theta_{g}$, with $\theta_{g}\in[0,1]$; the limit cases correspond, respectively, to a uniform temperature $T_{0}$, and a temperature gradient from 00 K to $2T_{0}$; the median temperature is $T_{0}$ in all situations.
@@
Steady-state characteristics Chemical force $\nicefrac{{\mathcal{A}}}{{T}}$ J.mol-1.K-1 $\alpha$ $=\nicefrac{{\mathcal{A}}}{{RT}}=\ln\nicefrac{{K}}{{Q}}$
@@
Table 1: Dimensionalized and nondimensionalized parameters, classified as characteristic dimensions, chemical characteristics, primary parameters (control parameters for the description of a given system), secondary parameters (used for the simplification of the final equations), and steady-state characteristics (describing the resulting steady dynamic properties of the system). $R=8.314$ J.mol-1.K-1.
@@
All kinetic and thermodynamic constants can be expressed as a function of these parameters at any point in the system, based on the Arrhenius equation (see Eq. (25aea)-(25aec)). $\kappa$ and $\lambda$ can be combined into two secondary parameters $\kappa^{\prime}$ and $\lambda^{\prime}$, leading to simpler mathematical representations. They represent the reaction and exchange asymmetry, respectively, with a maximal value of $1$ in the case of a perfectly symmetric process, and 00 for a process totally displaced in a given direction.
The resulting system behavior in the steady-state was then characterized in terms of three dimensionless parameters: chemical force $\alpha$, chemical flux $\varphi$, and entropic dissipation $\sigma$, which can be expressed as:
@@
II.2.2 Establishment of a nonequilibrium steady-state
The difference in temperature implies a difference in the equilibrium constant in each compartment. Thus, a temperature gradient necessarily leads to a frustrated state which is characterized by the impossibility of simultaneously reaching chemical equilibrium in each compartment (which would imply that each compound coexists at different concentrations in both compartment), and the transport equilibrium between each compartment (which would imply identical concentrations for each compound in both compartments).
This results in a nonequilibrium state, whose distance from the equilibrium can be quantified by the chemical force $\alpha$, as the sum of the chemical forces of each chemical reaction in each compartment and the chemical forces of each chemical exchange between the two compartments. It can then be expressed for low values of $\theta_{g}$ (see Eqs. (25aiar)-(25aias)) as:
@@ II.2.2 Establishment of a nonequilibrium steady-state
In all cases, a steady chemical force is thus sustained by the temperature gradient. It is proportional to the gradient intensity $\theta_{g}$ and the reaction enthalpy $\eta$, and does not depend on the other parameters.
In response to this nonzero chemical force, a circular reaction-diffusion flux is established. Chemical reaction fluxes are processed in opposite directions in each compartment, compensated by the continuous exchange of each compound in the opposite direction (see Fig. 2A). By analogy with electric circuits $\alpha$ can be interpreted as the potential difference, and $\varphi$ as the intensity. These are expressed as (see Eqs. (25aiba) and (25aibd)):
@@ II.2.3 Process Distribution
This nonequilibrium steady reaction-diffusion flux leads to a continuous dissipative process characterized by $\sigma$, as an analog to electric power dissipation[36] according to Eq. 2. The energy extracted from the temperature gradient is dissipated differently in each process (see Fig. 2B), and can be decomposed as a sum of terms specific to each process, according to each chemical resistance as (see Eqs. (25aian)-(25aiap)):
@@
  • If $\lambda^{\prime}\delta\ll\rho$, then $\sigma\approx\sigma_{d}$; this is the case for a fast reacting system. Chemical equilibrium is observed in each separate compartment, and most of the dissipative process is directed to the chemical exchanges between compartments.
In the intermediate regime $\rho\sim\lambda^{\prime}\delta$, both chemical reaction and transport are actively processed.
@@
Qualitative changes are observed as functions of $\varepsilon\theta_{g}$. When $\varepsilon\theta_{g}\ll 1$, the dissipation due to the chemical reaction is equally shared between both compartments as $\sigma_{ra}\approx\sigma_{rb}$. This dissipation is asymmetric for higher $\varepsilon\theta_{g}$ values, increasing the dissipation in the colder compartment, and reducing it by the same value in the warmer compartment.
This asymmetry originates from the slow down of the chemical reaction in the cold compartment. The chemical flux $\varphi$ is necessarily of the same intensity in each compartment to guarantee a steady-state; this implies an increase in the chemical force $\alpha$ in the cold compartment owing to a less efficient chemical relaxation towards equilibrium, and its decrease in the warm compartment owing to a more efficient relaxation. Consequently, the more dissipative compartment is necessarily the colder one.
@@
It is symmetrically distributed between the two compounds when $\lambda\kappa=1$. A dissymetrization of this exchange is quantified by $\lambda\kappa$, with a $\sigma_{d}\approx\sigma_{d1}$ for $\lambda\kappa\gg 1$ (the exchange dissipation is then governed by $U_{1}$), and with a $\sigma_{d}\approx\sigma_{d2}$ for $\lambda\kappa\ll 1$ (the exchange dissipation is then governed by $U_{2}$).
@@ II.2.4 Optimal conditions
To maximize the efficiency of the nonequilibrium process, it is first necessary to maximize its distance from the equilibrium, i.e. to maximize $\alpha$. According to Eq. (3), this implies to maximize both $\theta_{g}$ and $\eta$. Thus, high temperature gradients coupled with a high-enthalpy chemical reaction should be favored as a necessary condition.
@@
Overall, this implies that each process should be fast and symmetrical at mean temperature $T_{0}$. The corresponding simultaneous optimization of $\alpha$ and $\rho$ implies the maximization of the dissipated entropy $\sigma$, and thus of the energy extracted from the temperature gradient.
The optimization can also be based on the efficiency of the global system in extracting and using energy from the temperature gradient, as assessed by the entropic yield (Eqs. (25aibc)). The majority of the entropy is dissipated for matter transport between the compartments for $\lambda^{\prime}\delta>\rho$, and for the nonequilibrium reaction inside each compartment for $\lambda^{\prime}\delta<\rho$.
@@
Consequently, general optimization involves maximizing $\theta_{g}$, $\eta$, $\lambda^{\prime}$ and $\kappa^{\prime}$. Additionally, $\lambda^{\prime}\delta$ should be of the same order of magnitude as $\rho$, with $\lambda^{\prime}\delta>\rho$ for sustaining chemical reactions, $\lambda^{\prime}\delta<\rho$ for sustaining chemical transport, and $\lambda^{\prime}\delta\simeq\rho$ for sustaining reaction-diffusion cycles.

### v2

# spacing=nonfrench

@@ Abstract
We present a theoretical study of 1D chemical systems subjected to temperature gradients. A complete theoretical framework describes the behavior of the system induced by various temperature profiles. An exact mathematical derivation was established for a simple two-compartment model and was generalized to arbitrary reaction-diffusion systems based on numerical models. An experimental system was eventually scaled and tuned to optimize either nonequilibrium chemical transport or reaction.
The relevant parameters for this description were identified; they focused on the system symmetry for chemical reaction and transport. Nonequilibrium thermodynamic approaches lead to a description analogous to electric circuits. Temperature gradients lead to the onset of a steady chemical force, which maintains steady reaction-diffusion fluxes moderated by chemical resistance. The system activity was then assessed using the entropy production rate as a measure of its dissipated power.
The chemical characteristics of the system can be tuned for general optimization of the nonequilibrium state or for the specific optimization of either transport or reaction processes. The shape of the temperature gradient can be tailored to precisely control the spatial localization of active processes, targeting either precise spatial localization or propagation over large areas. The resulting temperature-driven chemical system can in turn be used to drive secondary processes into either nonequilibrium reaction fluxes or concentration gradients.
Free energy transduction, coupled reactions, reaction-diffusion, steady state, temperature gradient, thermodynamics of nonequilibrium, chemical fluxes, entropy rate.
preprint: Submitted to Phys Rev E
@@ I Introduction
Thermodynamics fixes the final composition in which a reactive system reaches equilibrium under chemical and physical constraints. In practice, it governs the realization of multiple chemical reactions of academic and industrial interest, as well as the elementary steps of most separation processes, such as distillation or chromatography. In contrast, kinetics drives the trajectory of the evolution of a reactive system. It also governs the dynamic behavior of the latter when sustained in an nonequilibrium steady state, which can exhibit much richer phenomena (e.g. oscillations, waves, patterns, and chaos) than in the equilibrium state. In particular, such nonequilibrium steady states are encountered in living biological matter, where they control its dynamics[1, 2] and lead to the propagation of a sustained energy flux throughout metabolism by exploiting free energy transduction[3, 4, 5].
@@
The dynamic behavior of the nonequilibrium steady states is theoretically understood[6, 7, 8]. In contrast, the issue of steadily sustaining an nonequilibrium reactive state has received much less attention and remains a major challenge[9]. In practice, its realization requires either introducing constraints to prevent the effective relaxation of the reactive processes towards equilibrium[10, 11, 12], or steadily applying a periodic excitation on the reactive system.[13] We previously demonstrated the relevance of the first approach by frustrating the relaxation of reactive processes by the interference of molecular motion[14, 15], a strategy to generate reaction-diffusion cycles that are thought to establish spatial gradients of signaling activities in living cells[16, 17]. However, the efficiency of this first approach has not been theoretically analyzed, and is the object of the present manuscript. xpired. In our previous work[14, 15], we used light as an energy source to maintain a reactive system in a nonequilibrium state. In the present study, we favored a more general approach and adopted the application of a temperature gradient. Despite its chemical relevance, the behavior of reactive media in contact with two thermostats at different temperatures has included glorious names (including Nernst[18], Dirac[19], Prigogine[20, 21]). Several papers have been published since then, but have been mainly concerned with the gas phase and the question of how a chemical reaction that occurs in a system modifies its thermal conductivity[22, 23, 24, 25, 26, 27, 28, 29, 30]. Recently, heat flow has been applied to generate proton gradients and pH oscillations in microscale aqueous solutions[31, 32]. In the latter works, the reactive system was often assumed to be at chemical equilibrium within the temperature gradient.
@@
In this paper, we level off this assumption, which has been predicted to generate attractive behavior such as dissipation-driven selection of states in nonequilibrium chemical networks[33], or emergent thermophoretic behavior in chemical reaction systems[34]. More specifically, we identify the optimal regime for shifting away a reactive system from chemical equilibrium at a steady state. We analyze optimization of the flux of the reaction and diffusion processes and of the corresponding entropy production[35], which are key elements in the propagation of energy into a set of coupled reactions. To achieve this optimization, we benefit from a fruitful analogy between entropy production by a chemical reaction and the ohm dissipated power by electrical resistors. We then identify the conditions for the spatial localization or delocalization of nonequilibrium activity. Finally, we assessed the possibility of propagating this primary energy transfer from the heat sources and sinks to secondary athermic reaction or transport processes.
@@ II.1 Chemical model
This represents a generic reaction, in its simplest form. It can be easily extended to more complex mechanisms, as detailed further, namely for the description of second-order reactions or more complex mechanisms.
@@
Figure 1: Description of the chemical system. A, general representation; B, two-compartment model; C, unstirred model; D-F, applied temperature gradient profiles.
@@ II.2 Chemical response to a temperature gradient between two homogeneous compartments
The general behavior of a reaction driven by temperature gradients was first characterized in the limits of two compartments, each of which was homogeneous in temperature and concentration. This corresponds to two stirred systems; each compartment $A/B$ is connected to the other by chemical exchange of the reaction rate constants $\delta_{i}$ for each compound of concentration $u_{i}$.
This setup can be reduced to a set of ordinary differential equations with a set of parameters for each compartment. Analytical solutions for the resulting steady state were obtained. A second-order Taylor development as a function of the temperature gradient $\theta_{g}$ leads to simpler but more informative solutions.
@@ II.2.1 Parameters
Identifying the relevant parameters for the description of the system is a critical step. We focused on describing the system in general terms by assessing the global symmetry or asymmetry of the parameters, as this leads to a description of the system behavior in the simplest mathematical terms.
The general properties of the system are described based on its characteristic dimensions: concentration $c_{0}$ (corresponding to the total concentration), temperature $T_{0}$ (corresponding to the median system temperature) and time $t_{0}$. All subsequent system parameters were nondimensionalized using these characteristic parameters (see Table 1).
@@
Steady state characteristics Chemical force $\nicefrac{{\mathcal{A}}}{{T}}$ J.mol-1.K-1 $\alpha$ $=\nicefrac{{\mathcal{A}}}{{RT}}=\ln\nicefrac{{K}}{{Q}}$
@@
Table 1: Dimensionalized and nondimensionalized parameters, classified as characteristic dimensions, chemical characteristics, primary parameters (control parameters for the description of a given system), secondary parameters (used for the simplification of the final equations), and steady state characteristics (describing the resulting steady dynamic properties of the system). $R=8.314$ J.mol-1.K-1.
The system is thus described in the nondimensionalized concentrations $u_{i,j}$ of compound $i$ in compartment $j$, relative temperature deviation $\theta$, and time $\bar{t}$. Each compartment is thermostated at temperature $\theta_{a}=-\theta_{g}$ and $\theta_{b}=+\theta_{g}$, with $\theta_{g}\in[0,1]$; the limit cases correspond, respectively, to a uniform temperature $T_{0}$ and a temperature gradient from 00 K to $2T_{0}$; the median temperature is $T_{0}$ in all situations.
@@
All kinetic and thermodynamic constants can be expressed as a function of these parameters at any point in the system, based on the Arrhenius equation (see Eqs. (25aea)-(25aec)). $\kappa$ and $\lambda$ can be combined into two secondary parameters $\kappa^{\prime}$ and $\lambda^{\prime}$, leading to simpler mathematical representations. They represent the reaction and exchange asymmetry, respectively, with a maximal value of $1$ for a perfectly symmetric process and 00 for a process totally displaced in a given direction.
The resulting system behavior in the steady state was then characterized in terms of three dimensionless parameters: chemical force $\alpha$, chemical flux $\varphi$, and entropic dissipation $\sigma$, which can be expressed as:
@@
II.2.2 Establishment of a nonequilibrium steady state
The difference in temperature implies a difference in the equilibrium constant in each compartment. Thus, a temperature gradient necessarily leads to a frustrated state, which is characterized by the impossibility of simultaneously reaching chemical equilibrium in each compartment (which would imply that each compound coexists at different concentrations in both compartments) and the transport equilibrium between each compartment (which would imply identical concentrations for each compound in both compartments).
This results in a nonequilibrium state, whose distance from equilibrium can be quantified by the chemical force $\alpha$, as the sum of the chemical forces of each chemical reaction in each compartment and the chemical forces of each chemical exchange between the two compartments. It can then be expressed for low values of $\theta_{g}$ (see Eqs. (25aiar)-(25aias)) as:
@@ II.2.2 Establishment of a nonequilibrium steady state
In all cases, a steady chemical force is thus sustained by the temperature gradient. It is proportional to the intensity of the gradient $\theta_{g}$ and the enthalpy of the reaction $\eta$ and does not depend on the other parameters.
In response to this non-zero chemical force, a circular reaction-diffusion flux is established. Chemical reaction fluxes are processed in opposite directions in each compartment, compensated by the continuous exchange of each compound in the opposite direction (see Fig. 2A). By analogy to electric circuits, $\alpha$ can be interpreted as the potential difference and $\varphi$ as the intensity. These are expressed as (see Eqs. (25aiba) and (25aibd)):
@@ II.2.3 Process Distribution
This steady nonequilibrium reaction-diffusion flux leads to a continuous dissipative process characterized by $\sigma$, as an analog of electric power dissipation[36] according to Eq. 2. The energy extracted from the temperature gradient is dissipated differently in each process (see Fig. 2B), and can be decomposed as a sum of terms specific to each process, according to each chemical resistance as (see Eqs. (25aian)-(25aiap)):
@@
  • If $\lambda^{\prime}\delta\ll\rho$, then $\sigma\approx\sigma_{d}$; this is the case for a fast-reaction system. Chemical equilibrium is observed in each separate compartment, and most of the dissipative process is directed to the chemical exchanges between compartments.
In the intermediate regime $\rho\sim\lambda^{\prime}\delta$, both the chemical reaction and the transport are actively processed.
@@
Qualitative changes are observed as functions of $\varepsilon\theta_{g}$. When $\varepsilon\theta_{g}\ll 1$, the dissipation due to the chemical reaction is equally shared between both compartments as $\sigma_{ra}\approx\sigma_{rb}$. This dissipation is asymmetric for higher values of $\varepsilon\theta_{g}$, increasing the dissipation in the colder compartment and reducing it by the same value in the warmer compartment.
This asymmetry originates from the slowing of the chemical reaction in the cold compartment. The chemical flux $\varphi$ is necessarily of the same intensity in each compartment to guarantee a steady state; this implies an increase in the chemical force $\alpha$ in the cold compartment due to a less efficient chemical relaxation toward equilibrium and its decrease in the warm compartment owing to a more efficient relaxation. Consequently, the more dissipative compartment is necessarily the colder.
@@
It is symmetrically distributed between the two compounds when $\lambda\kappa=1$. The asymmetrization of this exchange is quantified by $\lambda\kappa$, with a $\sigma_{d}\approx\sigma_{d1}$ for $\lambda\kappa\gg 1$ (the exchange dissipation is then governed by $U_{1}$), and with a $\sigma_{d}\approx\sigma_{d2}$ for $\lambda\kappa\ll 1$ (the exchange dissipation is then governed by $U_{2}$).
@@ II.2.4 Optimal conditions
To maximize the efficiency of the nonequilibrium process, it is first necessary to maximize its distance from the equilibrium, i.e., to maximize $\alpha$. According to Eq. (3), this implies maximization of both $\theta_{g}$ and $\eta$. Thus, high temperature gradients coupled with a high-enthalpy chemical reaction should be favored as a necessary condition.
@@
Overall, this implies that each process should be fast and symmetrical at mean temperature $T_{0}$. The corresponding simultaneous optimization of $\alpha$ and $\rho$ implies maximization of the dissipated entropy $\sigma$, and thus of the energy extracted from the temperature gradient.
Optimization can also be based on the efficiency of the global system in extracting and using energy from the temperature gradient, as assessed by the entropic yield (Eqs. (25aibc)). Most of the entropy is dissipated for the transport of matter between the compartments for $\lambda^{\prime}\delta>\rho$, and for the nonequilibrium reaction within each compartment for $\lambda^{\prime}\delta<\rho$.
@@
Consequently, general optimization involves maximizing $\theta_{g}$, $\eta$, $\lambda^{\prime}$, and $\kappa^{\prime}$. Furthermore, $\lambda^{\prime}\delta$ should be of the same order of magnitude as $\rho$, with $\lambda^{\prime}\delta>\rho$ to sustain chemical reactions, $\lambda^{\prime}\delta<\rho$ for sustaining chemical transport, and $\lambda^{\prime}\delta\simeq\rho$ for sustaining reaction-diffusion cycles.